### v1

@@ 3.4 Concept Isolation Constraints
While the Concept Injection Constraints effectively guarantee that objects will be placed within user-specified regions, it does not prevent the potential overlap or “infection” of customized concepts within these areas, as illustrated in Fig. 6(d). To preserve the integrity and distinctiveness of each concept within its designated region, we introduce Concept Isolation Constraints. This approach is divided into two main components: Concept Region Mask and Region Perceptual Restriction. Both elements are integrated within the self-attention mechanism of the U-Net block, ensuring that each concept remains isolated and unaffected by others, thereby maintaining the purity of concepts in their target regions.
Concept Region Mask: The self-attention mechanism creates connections among all query elements, essential for maintaining the distinct characteristics of each concept. To preserve the distinctiveness of each concept, we adopt the concept region mask strategy, guided by a given layout condition as depicted in Fig. 3(b). This methodology limits the interaction between queries within a specific concept region and those in other concept regions, as demonstrated in Fig. 3(c). Consequently, this ensures the preservation of each concept’s unique characteristics, free from the influence of neighboring concepts.
Region Perceptual Restriction: Due to down-sampling and residual convolution operations in the Stable Diffusion framework, concept features might spread into the elements designated for background areas, as highlighted by the yellow square in Fig. 3(b). To mitigate the risk of concept feature leakage into unintended regions, we employ Region Perceptual Restriction, aimed at minimizing interaction between queries of the foreground and background areas. This technique ensures that each concept remains distinct and unaffected by the features of the background feature, thereby preserving the uniqueness and integrity of each concept within the synthesized image. This formulated as
$\mathcal{L}_{region}=\sum_{n=1}^{n}(\frac{1}{P}\sum^{P}_{P=0}\textbf{topk}(\bar{A}_{[M_{n},\mathbf{1}-{M}_{n}]},P)),$ (8)
where the $\bar{A}_{[M_{n},\mathbf{1}-{M}_{n}]}$ refers to the self-attention map obtained through a matrix slicing operation across the channel dimension.
@@
$\mathcal{L}=\mathcal{L}_{ce}+\alpha\mathcal{L}_{fill}+\beta\mathcal{L}_{region},$ (9)
@@
Following Boxdiff[39], the step size $\phi_{t}$ decays linearly with each timestep. By incorporating the previously mentioned constraints, $z^{\prime}_{t}$ is directed at each timestep to foster the generation of customized concepts within designated locations, while preventing the leakage of concept features into areas associated with other concepts. Subsequently, $z^{\prime}_{t}$ is utilized as the input for the U-Net for the ensuing inference step $z^{\prime}_{t}\xrightarrow{U-Net}z_{t-1}$. This strategic guidance ensures the precise synthesis of target objects within the user-specified layout regions.
@@ 3.5 Latent Re-initialization
We discovered that traditional LoRA is not ideally suited for generating specific local areas, because it is trained on the generation of entire images. This discrepancy between training and inference processes can result in misalignment. To address this issue, we propose re-initializing the latent space to better accommodate the integration of concept-specific LoRAs.
Before the denoising phase, we re-initialize the latent space to better integrate concept-specific LoRAs. This begins with sampling a random latent vector from a Gaussian distribution, then applying the LoRA-Composer process, which performs one-step update using Eq. 10. Afterward, the cross-attention map $A_{n}^{t}$ is generated, based on the latent query mapping and the textual embedding mapping for each local prompt. The subsequent step focuses on identifying the area with the highest scores to replace the predefined concept mask area $M_{n}$. Specifically, the indices of area $\hat{i_{n}},\hat{j_{n}}$ of certain concept is determined as follows:
$\hat{i}_{n},\hat{j}_{n}=\underset{i\in 0,1\dots w,j\in 0,1\dots h}{\mathrm{arg\,max}}\ \sum_{i}^{i+\mathbb{W}}\sum_{j}^{j+\mathbb{H}}\textbf{crop}(A^{c}_{n},[i,j],\mathbb{W},\mathbb{H}).$ (11)
In this context, $\mathbb{W}$ and $\mathbb{H}$ represent the width and height of the concept mask $M_{n}$, respectively. The variables $w$ and $h$ denote the width and height of the cross-attention map $A^{c}_{n}$. The function $\textbf{crop}(A,[i,j],\mathbb{W},\mathbb{H})$ refers to cropping the attention map to a shape of $\mathbb{W},\mathbb{H}$ with the top-left coordinate at $[i,j]$. Finally, the latent is normalized to a standard Gaussian distribution to prepare for subsequent processing steps.
Figure 4: Three highlights of LoRA-Composer: a) full image customization in multi-style; b) manipulating interactions and attributes; c) multi-condition control generation.
@@ 4.1 Experimental Setup
Datasets. For a thorough evaluation of LoRA-Composer, follow Mix-of-Show[7], we compile a dataset featuring characters, animals, and backgrounds in both realistic and anime styles, encompassing a total of 16 customized subjects. Through comprehensive experimentation across diverse subject combinations, we demonstrate the superior performance of our approach.
Evaluation metrics. Our evaluation of the LoRA-Composer approach utilizes two metrics for customized generation, as proposed in prior studies [18, 7, 13]. (1) Image similarity: This metric assesses the visual resemblance between the generated images and the target subjects in the CLIP [24] image embedding. In our multi-concept generation task, we compute the image similarity for the generated images about each target concept individually and then determine the average value. (2) Textual similarity: This metric evaluates the mean similarity between generated images and their associated textual prompts, utilizing the CLIP model [24] for quantification. For our multi-concept generation task, for the foreground concepts, we first isolate each concept within the generated images. Subsequently, we extract every concept token $V$ in local prompt $P^{n}_{L}$ and substitute it with the concept’s class name before computing the textual similarity. When evaluating background concepts, we input the entire image. The final score is the average of these similarity values.
Baseline. To assess the quality of our generated images, we compare our approach against four leading competitors in the field. Cones2[18] leverages text embeddings to support arbitrary combinations of concepts without necessitating model refinement. Mix-of-Show[7] employs gradient fusion to integrate multiple concepts into a base model seamlessly. Both Anydoor and Paint by Example facilitate multi-concept generation by utilizing networks trained specifically for inpainting tasks. More implementation details are shown in the Appendix.
@@ 4.2 Visualization Results
Our methodology enables extensive customization of the entire image, addressing both foreground and background components and accommodating a wide range of styles, from anime to realistic. This broad customization capability is illustrated in Fig. 4(a). showcase our approach’s versatility in adapting to artistic expressions. Additionally, our method enables precise manipulation of interactions and attributes, such as shaking hands, wearing hats, and holding teddy bears in the picture, directly through textual prompts. This capability is showcased in Fig. 4(b). Moreover, our framework is designed for flexibility, capable of generating images under multiple conditions. It adeptly integrates specific constraints such as edge detection or pose estimation to guide the image synthesis process. This capacity for accommodating additional image-based conditions, as detailed in Fig. 4(c), highlights the adaptability of our approach in meeting varied and complex generation requirements.
@@
Table 1: LoRA-Composer’s performance in generating anime and realistic style concepts is benchmarked against baseline models. T refers to textual similarity, I refers to image similarity, with an asterisk * indicating the use of image-based conditions. The highest scores in each column are marked in bold for clarity.
@@ 4.3 Quantitative Results
As detailed in Tab. 1, our LoRA-Composer surpasses prior techniques in image similarity, showcasing its effectiveness across both anime and realistic styles. Conversely, inpainting-based methods such as Anydoor and Paint by Example exhibit higher text similarity. This is because these methods specialize in inserting subjects into user-defined locations through reference images, focusing more on aligning with textual descriptions. Our method achieves significant improvements over Cones2, thanks to the efficacy of our LoRA-Composer block. These data illustrate that our method excel in precisely drawing out target concepts.
@@ 4.4 Qualitative Comparison
LoRA-Composer is evaluated without using image-based conditions across four benchmarks in multi-concept customization scenarios. The results are displayed in Fig. 5. It can be seen that Cones2[18] faces concept confusion, unable to clearly capture concept characteristics. In contrast, inpainting-based models like Anydoor[4] and Paint by Example[40] retain certain concept features, such as glasses and gray curly hair as illustrated in the first row, but struggle with detailed attributes, especially facial features. Mix-of-Show[7] also suffers from concept confusion, as seen in the first row where glasses are incorrectly placed on a man. Furthermore, in the second and third rows, two people are completely vanishing. Differently, our method successfully synthesizes images that accurately incorporate all subjects with their correct characteristics, showcasing enhanced performance in multi-concept synthesis and attribute accuracy.
Figure 5: Qualitative comparison with baselines. For each case, we use the same seeds across all of the methods.
@@
4.5 Ablation Study
To demonstrate the efficacy of each component in our method, we choose a challenging scenario with five and four concepts for this section. As illustrated in Fig. 6(c), the positions of the anime girl and the woman within the red box diverge from the layout conditions outlined in Fig. 6(a). This discrepancy serves as evidence that omitting Latent Re-Initialization (LR) hampers the precise placement of concepts, mainly because of the lack of spatial priors. Subsequently, as shown in Fig. 6(d), eliminating the Concept Isolation Constraints (CI) results in the blending of concept characteristics, such as the confusion observed in the anime girl’s haircut and the distortion of the man’s hair color. Without CI, concepts begin to overlap and influence each other, resulting in a disruption of harmony and coherence in the overall image composition. Finally, as shown in Fig. 6(e), eliminating the Concept Enhancement Constraints (CE) results in the disappearance of concepts. However, thanks to the presence of Region-Aware LoRA Injection, the system retains the capability to insert concepts, albeit with reduced precision in their placement and representation. This highlights each element’s critical role in achieving precise and harmonious concept integration.
To substantiate our findings as non-coincidental, we conducted a comprehensive quantitative evaluation. As detailed in Tab. 2, our analysis demonstrates the individual and collective impacts of Concept Enhancement Constraints (CE), Latent Re-Initialization (LR), and Concept Isolation Constraints (CI) on model performance. CE lead to significant improvements across all performance metrics, showcasing its effectiveness in activating concepts, evidenced by an increase of around 0.1 in mean image similarity. LR further contributed to these enhancements by refining region-specific priors. CI played a crucial role in preserving the distinctiveness of concept traits and enhancing model robustness.
Figure 6: Visualized results from ablation study on individual components. “L” stands for Latent Re-initialization, “CI” denotes Concept Isolation Constraints, and “CE” signifies Concept Enhancement Constraints within Concept Injection Constraints.
@@ 5 Conclusion
In this paper, we introduce LoRA-Composer, a novel approach designed to seamlessly integrate multiple concepts within a single image. We explore two prevalent issues in multi-concept customization: concept vanishing and concept confusion. To this end, we employ Concept Injection Constraints to combat concept vanishing, while Concept Isolation Constraints and Latent Re-initialization to alleviate concept confusion. Our experiments highlight the capability of LoRA-Composer to customize entire images, including both background and foreground elements, and to manipulate the interactions and attributes of various concepts through textual prompts. Compared to traditional methods, LoRA-Composer offers enhanced flexibility and usability, allowing users to generate images with fewer conditions and readily accessible LoRA techniques. Furthermore, we demonstrate the method’s ability to achieve high-fidelity combinations of multiple concepts, underscoring its practical utility in complex image-generation tasks.
Supplementary Material
Considering the space limitation of the main text, we provided more results and discussion in this supplementary material, which is organized as follows:
  • Section 0.A: implementation details of our approach and the baseline models.
  • Section 0.B: more detailed experiments analysis and discussion.
    • Section 0.A.1: our default setting in experiments and ablation study.
    • Section 0.B.1: ablation study on Concept Enhancement Constraints (in Eq. 7) and Concept Isolation Constraints (in Eq. 8).
    • Section 0.B.2: comparison with Mix-of-Show, under the image-based conditions.
    • Section 0.B.3: assess the human preference between our method and baseline approaches.
    • Section 0.B.4: more visual results of LoRA-Composer.
  • Section 0.C: discussion of our method’s potential negative society impact.
  • Section 0.D: failure cases and discussion.
Appendix 0.A Implementation Detail
@@
Sample Details. For all experiments and evaluations in this paper, we use the DPM-Solver[19], implementing adaptive sampling steps to enhance computational efficiency. Specifically, if the loss (as described in Eq. 9) ceases to decrease, we stop the process, thereby accelerating the overall procedure. For Eq. 9, the relative coefficients are set as $\alpha=0.25$ and $\beta=0.8$.
0.A.1 Evaluation Setting
Baseline Implementation Detail. For Cones2[18], we utilize the official implementation provided at the repository [^3] https://github.com/ali-vilab/Cones-V2. The specified training configurations are a batch size of 4, a learning rate of 5e-6, and a total of 4000 training steps. To ensure consistency across experiments, we employ the same seed for image generation. Given that Paint by Example[40] [^4] https://github.com/Fantasy-Studio/Paint-by-Example and Anydoor[4] [^5] https://github.com/ali-vilab/AnyDoor focus exclusively on real object inpainting, we ensure a fair comparison by limiting the comparison to real-world concepts. Specifically, our approach involves initially generating the background image using our model with the same prompt and seeds, while omitting the foreground prompts. Subsequently, their models are employed to introduce the foreground concepts. For Mix-of-Show[7], we utilize the same LoRAs for both real-world and anime concepts. We apply their gradient fusion technique [^6] https://github.com/TencentARC/Mix-of-Show/tree/research_branch separately to each concept type. Consistency across experiments is ensured by using the same seed for image generation, allowing for a direct comparison of outcomes.
Appendix 0.B Additional Experiments

### v2

@@ 3.3 Concept Isolation Constraints
While the concept injection constraints effectively guarantee that objects will be placed within user-specified regions, they cannot prevent the potential overlap or infection of customized concepts within these areas (see Fig. 6(d)). To preserve the integrity and distinctiveness of each concept within its designated region, we introduce concept isolation constraints. This approach is divided into two main components: concept region mask and region perceptual restriction. Both elements are integrated within the self-attention mechanism of the U-Net block, ensuring that each concept remains isolated and unaffected by others, thereby maintaining the purity of concepts in their target regions.
Concept Region Mask. The self-attention mechanism creates connections among all query elements, essential for maintaining the distinct characteristics of each concept. To preserve the distinctiveness of each concept, we adopt the concept region mask strategy, guided by a given layout condition as depicted in Fig. 3(b). This design limits the interaction between queries within a specific concept region and those in other concept regions, as demonstrated in Fig. 3(c). Consequently, it ensures the preservation of each concept’s characteristics, free from the influence of neighboring concepts.
Region Perceptual Restriction. Due to down-sampling and residual convolution operations in the U-Net framework, concept features might spread into the elements designated for background areas, as highlighted by the yellow square in Fig. 3(b). To mitigate the risk of concept feature leakage into unintended regions, we employ region perceptual restriction, aimed at minimizing interaction between queries of the foreground and background areas. This technique ensures that each concept remains distinct and unaffected by the features of the background feature, thereby preserving the uniqueness and integrity of each concept within the synthesized image. This formulated as
$\mathcal{L}_{r}=\frac{1}{S}\sum_{i=1}^{N}\textbf{topk}(\bar{A}{[M_{i},\mathbf{1}-{M}_{i}]},S),$ (7)
where the $\bar{A}{[M_{i},\mathbf{1}-{M}_{i}]}$ refers to the self-attention map obtained through a matrix slicing operation across the channel dimension.
@@
$\mathcal{L}=\mathcal{L}_{c}+\alpha\mathcal{L}_{f}+\beta\mathcal{L}_{r},$ (8)
@@
Following BoxDiff [21], the step size $\phi_{t}$ decays linearly with each timestep. By incorporating the previously mentioned constraints, $z^{\prime}_{t}$ is directed at each timestep to foster the generation of customized concepts within designated locations, while preventing the leakage of concept features into areas associated with other concepts. Subsequently, $z^{\prime}_{t}$ is utilized as the input for the U-Net for the ensuing inference step $z^{\prime}_{t}\xrightarrow{U-Net}z_{t-1}$. This strategic guidance ensures the precise synthesis of target concepts within the user-specified layout regions.
@@ 3.4 Latent Re-initialization
We discovered that traditional LoRA is not ideally suited for generating specific local areas, because it is trained without control in location. This discrepancy can result in imprecise locations for concept generation (see Fig. 6(c)). To address this issue, we propose re-initializing the latent space to better accommodate the integration of concept-specific LoRAs.
Our approach aims to identify the position within the latent space where the object is likely to appear and then align this position with the specified layout. Specifically, before the denoising phase, we initialize the latent space $z_{t}$ with Gaussian noise and apply the LoRA-Composer process for a one-step update using eq. 9. Afterward, a cross-attention map is generated based on the region latent query $Q_{i}$ and the textual embedding $K_{i}$ for each local prompt. We first compute each candidate area with the same shape as the layout. We then replace the layout area $z_{t}{[M_{i}]}$ with the highest scoring area in the candidate areas. Finally, the latent is normalized to a standard Gaussian distribution to prepare for subsequent processing steps. The aforementioned candidate areas can be expressed as:
$\hat{A_{i}}=\{\Phi(A_{i},[x,y],\mathbb{W},\mathbb{H})\},$ (10)
where $x\in\{0,1\dots w\},y\in\{0,1\dots h\}$. The $w$ and $h$ denote the width and height of the cross-attention map $A_{i}$. The function $\Phi(\cdot,[x,y],\mathbb{W},\mathbb{H})$ refers to cropping the attention map to a shape of $\mathbb{W},\mathbb{H}$ with the top-left coordinate at $[x,y]$. The variables $\mathbb{W}$ and $\mathbb{H}$ represent the width and height of the layout box.
Figure 4: Three highlights of LoRA-Composer, a) full image customization in multi-style; b) manipulating interactions and attributes; c) multi-condition control generation.
@@ 4.1 Experimental Setup
Datasets. For a thorough evaluation of LoRA-Composer, follow Mix-of-Show [2], we compile a dataset featuring characters, animals, and backgrounds in both realistic and anime styles, encompassing a total of 16 customized subjects (see Appendix Fig. 7). Through comprehensive experimentation across diverse subject combinations, we demonstrate the superior performance of our approach.
Evaluation metrics. Our evaluation of the LoRA-Composer approach utilizes two metrics for customized generation, as proposed in prior studies [9, 2, 3]. (1) Image similarity, which assesses the visual resemblance between the generated images and the target subjects in the CLIP [36] image embedding. (2) Textual similarity, which evaluates the similarity between generated images and their associated textual prompts, utilizing the CLIP model [36]. Please refer to Appendix B for details.
Baseline. We compare our approach against four leading competitors in the field. Cones2 [9] leverages text embeddings to support arbitrary combinations of concepts. Mix-of-Show [2] employs gradient fusion to integrate multiple concepts into a base model. Both Anydoor [7] and Paint by Example [6] facilitate multi-concept generation by utilizing networks trained specifically for inpainting tasks. More implementation details are shown in the Appendix B.
@@ 4.2 Visualization Results
Our broad customization capability is illustrated in Fig. 4(a), showcasing our approach’s versatility in adapting to a wide range of styles, from anime to realistic. Additionally, our method enables precise manipulation of interactions and attributes, such as shaking hands, wearing hats, and holding teddy bears in the picture, directly through textual prompts. This capability is showcased in Fig. 4(b). Moreover, our framework is designed for flexibility, capable of generating images under multiple conditions. It adeptly integrates specific constraints such as edge detection or pose estimation to guide the image synthesis process. This capacity for accommodating additional image-based conditions, as detailed in Fig. 4(c), highlights the adaptability of our approach in meeting varied and complex generation requirements. More visualization results are shown in Fig. 12 in the Appendix.
@@
Table 1: Quantitative comparison of LoRA-Composer with baselines in generating anime and realistic style concepts. T refers to textual similarity, I refers to image similarity, with an asterisk * indicating the use of image-based conditions. The highest scores in each column are marked in bold for clarity.
@@ 4.3 Quantitative Results
As detailed in Table 1, our LoRA-Composer surpasses prior methods in image similarity, showcasing its effectiveness across both anime and realistic styles. Conversely, inpainting-based methods such as Anydoor and Paint by Example exhibit higher text similarity. This is because these methods specialize in inserting subjects into user-defined locations through reference images, focusing more on aligning with textual descriptions. Our method achieves significant improvements over Cones2, thanks to the efficacy of our LoRA-Composer block.
@@
Figure 5: Qualitative comparison with baselines. For each case, we use the same seeds.
@@ 4.4 Qualitative Comparison
LoRA-Composer is evaluated without using image-based conditions across four benchmarks in multi-concept customization scenarios. The results are displayed in Fig. 5. It can be seen that Cones2 [9] faces concept confusion, unable to clearly capture concept characteristics. In contrast, inpainting-based models like Anydoor [7] and Paint by Example [6] retain certain concept features, such as glasses and gray curly hair as illustrated in the first row, but struggle with detailed attributes, especially facial features. These approaches also lead to disproportionate foreground elements and unnatural integration of foreground and background. Mix-of-Show [2] also suffers from concept confusion, as seen in the first row where glasses are incorrectly placed on a man to the left. Furthermore, in the second and third rows, two people are completely vanishing. Differently, our method successfully synthesizes images that accurately incorporate all subjects with their correct characteristics, showcasing enhanced performance in multi-concept synthesis and attribute accuracy. Additional qualitative comparison results and results using image-based condition guidance are provided Fig. 9 and Fig. 10 in the Appendix.
4.5 Ablation Study
To demonstrate the efficacy of each component in our method, we choose a challenging scenario with five and four concepts for this section. As illustrated in Fig. 6(c), the positions of the anime girl and the cat within the red box diverge from the layout conditions outlined in Fig. 6(a). This discrepancy serves as evidence that omitting latent re-initialization (LR) hampers the precise placement of concepts, mainly because of the lack of spatial priors. Subsequently, as shown in Fig. 6(d), removing the concept isolation constraints (CI) results in the blending of concept characteristics. This results in observable issues such as the confusion in the anime girl’s haircut, the distortion of the woman’s face, and the appearance of another cat in an incorrect area. Without CI, concepts begin to overlap and influence each other, resulting in a disruption of harmony and coherence in the overall image composition. Finally, as shown in Fig. 6(e), eliminating the concept enhancement constraints (CE) results in the disappearance of concepts. However, thanks to the presence of region-aware LoRA injection, the model retains the capability to insert concepts, though with diminished precision in their placement and representation. This highlights each element’s critical role in achieving precise and harmonious concept integration.
To substantiate our findings as non-coincidental, we conducted a comprehensive quantitative evaluation. As detailed in Table 2, our analysis demonstrates the individual and collective impacts of concept enhancement constraints (CE), latent re-initialization (LR), and concept isolation constraints (CI) on performance. CE lead to significant improvements across all performance metrics, showcasing its effectiveness in activating concepts, evidenced by an increase of around 0.1 in mean image similarity. LR further contributed to these enhancements by refining region-specific priors. CI played a crucial role in preserving the distinctiveness of concept traits and enhancing model robustness. More ablation results are shown in section C.2 in the Appendix.
@@ 4.5 Ablation Study
Figure 6: Visualized results from ablation study on individual components. “LR” stands for latent re-initialization, “CI” denotes concept isolation constraints, and “CE” signifies concept enhancement constraints within concept injection constraints.
@@ 5 Conclusion
In this paper, we introduce LoRA-Composer, a novel approach designed to seamlessly integrate multiple concepts within a single image. We explore two prevalent issues in multi-concept customization: concept vanishing and concept confusion. To this end, we employ concept injection constraints to combat concept vanishing, while concept isolation constraints alleviate concept confusion. Additionally, we propose latent re-initialization to provide precise region priors. Our experiments highlight the capability of LoRA-Composer to customize entire images, including both background and foreground elements, and to manipulate the interactions and attributes of various concepts through textual prompts. Compared to traditional methods, LoRA-Composer offers enhanced flexibility and usability, allowing users to generate images with fewer conditions and readily accessible LoRA techniques. Furthermore, we demonstrate the method’s ability to achieve high-fidelity combinations of multiple concepts, underscoring its practical utility in complex image-generation tasks.
Appendix A Preliminary
$\mathcal{L}_{sd}=\mathbb{E}_{z\sim\mathcal{E}(x),P,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau(P))\|_{2}^{2}\right],$ (11)
where $\epsilon_{\theta}$ is the denoising U-Net with learnable parameter $\theta$.
ED-LoRA [2] aims to augment the expressiveness of the embedding by employing a decomposed structure. ED-LoRA implements a layer-wise embedding strategy, following the method described in P+ [37], to forge a multi-faceted representation for the concept token ($V=[V_{rand},V_{class}]$). This involves adding a random variation ($V_{rand}$) and a class-specific component to the base embedding ($V_{class}$). Furthermore, it integrates a LoRA layer into the linear layers present in all attention modules of the text encoder and U-Net. This integration allows for a flexible adaptation of the model to specific concepts by modifying the linear layers in a low-rank manner, thereby enhancing the model’s ability to encode and synthesize images based on textual descriptions with high fidelity. We use it by default in all experiments.
Appendix B Implementation Detail
Sample Details. For all experiments and evaluations in this paper, we use the DPM-Solver [39], implementing adaptive sampling steps to enhance computational efficiency. Specifically, if the loss (as described in eq. 9) ceases to decrease, we stop the process, thereby accelerating the overall procedure. For this loss, the relative coefficients are set as $\alpha=0.25$ and $\beta=0.8$.
@@ Appendix B Implementation Detail
Evaluation metrics. In our multi-concept generation task, we crop each foreground concept within the generated images, while using the entire image for the background. We compute the image similarity for each target concept individually and then calculate the average value. In addition, we extract every concept token $V$ in local prompt $P^{i}$ and substitute it with the concept’s class name before computing the textual similarity. The final score is the average of these similarity values.
Baseline Implementation Detail. For Cones2 [9], we utilize the official implementation provided at the repository [^3] https://github.com/ali-vilab/Cones-V2. The training configurations specified include a batch size of 4, a learning rate of 5e-6, and a total of 4000 training steps. This setup requires approximately 10-15 minutes to execute on a single NVIDIA A100 GPU. To ensure consistency across experiments, we employ the same seed for image generation. Given that Paint by Example [6] [^4] https://github.com/Fantasy-Studio/Paint-by-Example and Anydoor [7] [^5] https://github.com/ali-vilab/AnyDoor focus exclusively on real object inpainting, we ensure a fair comparison by limiting the comparison to real-world concepts. Specifically, our approach involves initially generating the background image using our model with the same prompt and seeds, while omitting the foreground prompts. Subsequently, their models are employed to introduce the foreground concepts. For Mix-of-Show [2], we utilize the same LoRAs for both real-world and anime concepts. We apply their gradient fusion technique [^6] https://github.com/TencentARC/Mix-of-Show/tree/research_branch to integrate all of the LoRAs into the base model. Consistency across experiments is ensured by using the same seed for image generation, allowing for a direct comparison of outcomes.
Appendix C Additional Experiments
C.1 Default setting in experiments